From a boat on a lake, the angle of elevation to the top of a cliff is 21°33. If the base of the cliff is 373 feet From the boat, how high is the cliff(to the nearest foot)
Answer by ewatrrr(24785)

From a boat on a lake, the angle of elevation to the top of a cliff is 21°33.
If the base of the cliff is 373 feet From the boat, how high is the cliff(to the nearest foot)
tan = opp/adj
21°33 = 21 + 33/60 = 21.55°, radians = (21.55°/180°)pi = .376radians
373tan(.376radians) = 134ft rounded